# From attention to profit: quantitative trading strategy based on transformer

### v1

@@ 5.1 Overall performance
With the starting amount of RMB 5,000,000, the results of different strategies are obtained. The result of the experiment is shown in Table 2. For the monthly group, Month_1 shows an annual return of 17.35% and an annual excess return of 19.43%, indicating a robust performance over that period. Conversely, Month_3 shows a diminished annual return and annual excess return of 7.37% and 9.91% respectively. The turnover rate is notably high in the second month at 51.69%. The win rate ranges from 43.1% to 57.3%, with the first month showing the highest win rate. The Sharpe Ratio and Alpha, vary across the strategies, with some periods showing negative values, suggesting underperformance relative to the risk taken, while others show positive values.
@@
Strategy AR AER TR WR SR Alpha
Month_1 17.35% 19.43% 26.09% 57.8% 0.915 0.162
Month_2 9.91% 13.86% 51.69% 49.3% 0.289 0.102
Month_3 7.37% 9.91% 32.33% 51.6% 0.246 0.064
Week_1 -0.83% 1.31% 7.13% 46.4% -0.236 -0.03
Week_2 7.49% 10.81% 1.18% 49.2% 0.160 0.085
Week_3 12.3% 12.73% 1.39% 54.4% 0.372 0.116
Day_1 7.89% 11.4% 6.71% 43.1% 0.181 0.090
Day_2 10.23% 10.94% 6.51% 44.4% 0.279 0.097
Day_3 9.81% 10.03% 5.57% 44.6% 0.281 0.092
CSI300 1.77% \ \ \ \ \
@@
To demonstrate the advantages of the Transformer factor over traditional factors, 100 price-volume type factors from JoinQuant  [^3] https://www.joinquant.com/ (one of Chinese largest financial quantitative platforms) were selected for backtesting under the same trading strategy (Algorithm 1). Price-volume factors were used because these are calculated based on stocks’ prices, volumes, and turnover rates, which are similar to the training data used by Transformer factors (also used stock turnover rates and price data). The detailed performance of factors and benchmark is shown in Appendix A. On average, these factors achieved an annual return of -3.78%, an average excess return of -2.15%, and a Sharpe ratio of -0.36, where the benchmark return is -1.77%. The average maximum drawdown of the strategy was 44.88%, which means the potential worst-case scenario or the most extreme possible loss. The average volatility of the strategy ($\sigma_{p}$) was 0.26, with the benchmark at 0.197.
@@
Among these factor-based strategies, the Transformer factor performed well. Except for the win rate, which was slightly lower than a few strategies, it excelled in terms of annual return, annual excess return, $\sigma_{p}$, and Sharpe Ratio, ranking the best among the 101 factors. Figure 4 illustrates the return curves of the Transformer (Trans_Month_1) alongside the benchmark and some traditional factors. ”25_week_rank” represents the current price’s position over the past 25 weeks; ”EMAC20” and ”MAC20” are the 20-day index moving average and the stock’s 20-day moving average respectively; ”DAVOL20” and ”VOL20” represent the ratio of the 20-day average turnover rate to the 120-day average turnover rate, and the mean of the stock’s 20-day turnover rate; ”ROC20” is the price rate of change over 20 days; ”MAWVAD” is calculated as the product of the difference between the closing price and the opening price, divided by the range of the highest and lowest prices, all multiplied by the volume, accumulated over six days; ”Variance20” is the variance of the stock’s 20-day annualized returns. Most of these factors are compared based on a window of 20 timestamps, as the Transformer factor was also trained on 20 timestamps. In comparison with these factors, the Transformer factor (blue line) demonstrated a significantly better performance in terms of returns than the benchmark (orange line) and other factor strategies, expressing the improvement of the Transformer factor over traditional price-volume factors.
@@ 5.2 Half-life period test
To test the half-life period of the factor, another three portfolios based on $Month\_3$ were conducted. The first one is the stock pool used in the next month; the second is using it in the month after the month; the last one is using it in the third month after the last feature of inputs. For instance, in the first strategy, the trade will start in January 2020; in the second, it will start in February 2020; in the second, it will start in March 2020. For the period from January 2020 to February 2022, the strategy achieved a total return of 65.29% and an annual return of 17.87%. The excess return over this period was 66.39%, indicating a strong performance above a benchmark or risk-free rate. The Sharpe Ratio for this period was 0.863, which is a respectable figure that suggests a favorable risk-adjusted return. Additionally, the strategy generated an alpha of 0.164, demonstrating its ability to outperform the expected return on a risk-adjusted basis.
Table 3: Result of the same monthly strategy with different start month
Strategy Total return Annual return Excess return Sharpe ratio Alpha
2020/01-2022/02 65.29% 17.87% 66.39% 0.863 0.164
2020/02-2022/03 67.52% 18.21% 65.58% 0.940 0.161
2020/03-2022/04 56.41% 15.63% 52.96% 0.739 0.135
@@ 5.3 Discussion
The daily frequency strategies perform lower annual returns compared to others. This highlights the challenge of making profitable trades at higher frequencies - noise and volatility make reliable signals more difficult. The distribution of returns also shows increased dispersion at the daily level. However, daily strategies achieve more consistent turnover rates around 5-7%. This allows for portfolio adjustments while avoiding excessive trading costs. In contrast, monthly strategies see a turnover rate above 50% in some cases. There is likely an optimal point between 1 week to 1 month where matches market regularly.
@@ 6 Conclusion
This study proposed an approach to exploit transformer models for quantitative stock prediction and trading. We addressed the need for handling numerical input data rather than text and adapting the model for forecasting tasks rather than sequence-to-sequence problems common in NLP. To enable direct processing of numerical time series data, we replaced the word embedding layer with a standard linear layer and removed the output masking operations. We also simplified the decoder to produce a probability distribution over future price movements rather than autoregressively generating token sequences.
@@
There is further scope to enhance the framework, for instance by incorporating additional signals like news and fundamentals as auxiliary inputs. The self-attention mechanism could likely be improved to encode greater temporal context. Or use transformer-based models such as GPT-4 [45] and Claude3 [46] to establish the strategy as these models are fine-tuned and may perform better in quantitative finance tasks.
Overall, our work illustrates the viability of transformers for financial data modeling, provides a flexible framework to model the market from multiple perspectives of financial time series and develops profitable trading strategies that consider different variables. With further research to solve the above limitations, we believe such ML-based quantitative methods hold rich potential. The framework provides a flexible foundation for better understanding markets and developing profitable trading strategies.
@@ 7 Acknowledgements and declaration
Appendix A Detailed backtest results between 100 factors and the transformer factor
All the data of factors comes from JoinQuant [^4] https://www.joinquant.com/, which is one of the most common quantitative finance platforms. Detailed descriptions of the following factors are at JoinQuant Factors.
Seven metrics are used to evaluate the performance of each factor:
  • [label=•]
@@ Appendix A Detailed backtest results between 100 factors and the transformer factor
  • Sharpe Ratio (SR) is a measure of risk-adjusted return that can describe the additional earnings an investor receives per unit of increase, which is shown in Eq.(8).
@@
  • The Alpha measures the portfolio in generating returns above the market benchmark. A positive alpha indicates that the investment portfolio has achieved a higher return than its benchmark after risk adjustment.
@@
  • Beta is the correlation between the portfolio and systemic risk, reflecting the sensitivity of the strategy to the change in the market.

### v2

@@ 5.1 Overall performance
The result of the experiment is shown in Table 2. For the monthly group, Month_1 shows an annual return of 17.35% and an annual excess return of 19.43%, indicating a robust performance over that period. Conversely, Month_3 shows a diminished annual return and annual excess return of 7.37% and 9.91% respectively. The turnover rate is notably high in the second month at 51.69%. The win rate ranges from 43.1% to 57.3%, with the first month showing the highest win rate. The Sharpe Ratio and Alpha, vary across the strategies, with some periods showing negative values, suggesting underperformance relative to the risk taken, while others show positive values. In the comparison of 99% VaR, the strategies in monthly frequency perform better than others, where Month_1 and Month_3 are lower than 3%, which means the portfolio maximum loss 2.81% and 2.3% of return a day in 99% confidence.
@@
Strategy AR AER TR WR SR Alpha VaR
Month_1 17.35% 19.43% 26.09% 57.8% 0.915 0.162 2.81
Month_2  9.91% 13.86% 51.69% 49.3% 0.289 0.102 3.61
Month_3  7.37%  9.91% 32.33% 51.6% 0.246 0.064 2.3
Week_1 -0.83%  1.31% 7.13% 46.4% -0.236 -0.030 3.05
Week_2  7.49% 10.81% 1.18% 49.2% 0.160 0.085 3.73
Week_3 12.3% 12.73% 1.39% 54.4% 0.372 0.116 3.77
Day_1  7.89% 11.4% 6.71% 43.1% 0.181 0.090 3.92
Day_2 10.23% 10.94% 6.51% 44.4% 0.279 0.097 3.91
Day_3  9.81% 10.03% 5.57% 44.6% 0.281 0.092 4.02
CSI300 1.77% \ \ \ -0.015 \ 3.19%
@@
To demonstrate the advantages of the quantformer factor over traditional factors, 100 price-volume type factors from JoinQuant  [^3] https://www.joinquant.com/ (one of China’s largest financial quantitative platforms) were selected for backtesting under the same trading strategy (Algorithm 1). Price-volume factors were used because these are calculated based on stocks’ prices, volumes, and turnover rates, which are similar to the training data used by quantformer factors (also used stock turnover rates and price data). The detailed performance of factors and benchmark is shown in Appendix A. On average, these factors achieved an annual return of -3.78%, an average excess return of -2.15%, and a Sharpe ratio of -0.36, where the benchmark return is -1.77%. The average maximum drawdown of the strategy was 44.88%, which means the potential worst-case scenario or the most extreme possible loss. The average volatility of the strategy ($\sigma_{p}$) was 0.26, with the benchmark at 0.197.
@@
Among these factor-based strategies, the quantformer factor performed well. Except for the win rate, which was slightly lower than a few strategies, it excelled in terms of annual return, annual excess return, $\sigma_{p}$, Sharpe Ratio and sortino ratio ranking the best among the 101 factors. The 99% VaR of quantformer factor is in the quantile of the first 10% among all factors. Figure 5 illustrates the return curves of the quantformer (QF_Month_1) alongside the benchmark and some traditional factors. “25_week_rank” represents the current price’s position over the past 25 weeks; “EMAC20” and “MAC20” are the 20-day index moving average and the stock’s 20-day moving average respectively; “DAVOL20” and “VOL20” represent the ratio of the 20-day average turnover rate to the 120-day average turnover rate, and the mean of the stock’s 20-day turnover rate; “ROC20” is the price rate of change over 20 days; “MAWVAD” is calculated as the product of the difference between the closing price and the opening price, divided by the range of the highest and lowest prices, all multiplied by the volume, accumulated over six days; “Variance20” is the variance of the stock’s 20-day annualized returns. Most of these factors are compared based on a window of 20 timestamps, as the quantformer factor was also trained on 20 timestamps. In comparison with these factors, the quantformer factor (blue line) demonstrated a significantly better performance in terms of returns than the benchmark (orange line) and other factor strategies, expressing the improvement of the quantformer factor over traditional price-volume factors.
@@ 5.2 Half-life period test
To test the half-life period of the factor, another three portfolios based on $Month\_3$ were conducted. The first one is the stock pool used in the next month; the second is using it in the month after the month; the last one is using it in the third month after the last feature of inputs. For instance, in the first strategy, the trade will start in January 2020; in the second, it will start in February 2020; in the second, it will start in March 2020. For the period from January 2020 to February 2022, the strategy achieved a total return (TotalR) of 65.29% and an annual return of 17.87%. The total excess return (TER) over this period was 66.39%, indicating a strong performance above a benchmark or risk-free rate. The Sharpe Ratio for this period was 0.863, which is a respectable figure that suggests a favorable risk-adjusted return. Additionally, the strategy generated an alpha of 0.164, demonstrating its ability to outperform the expected return on a risk-adjusted basis.
Table 3: Result of the same strategy with different starting month
Strategy TotalR AR TER SR Alpha CSI300
2020/01-2022/02 65.29% 17.87% 66.39% 0.863 0.164 -0.66%
2020/02-2022/03 67.52% 18.21% 65.58% 0.940 0.161 1.17%
2020/03-2022/04 56.41% 15.63% 52.96% 0.739 0.135 2.26%
@@
The daily frequency strategies perform lower annual returns and higher risk compared to others This highlights the challenge of making profitable trades at higher frequencies - noise and volatility make reliable signals more difficult. The distribution of returns also shows increased dispersion at the daily level. However, daily strategies achieve more consistent turnover rates around 5-7%. This allows for portfolio adjustments while avoiding excessive trading costs. In contrast, monthly strategies see a turnover rate above 50% in some cases. There is likely an optimal point between 1 week to 1 month where matches market regularly.
5.3 Training under different scales
Understanding data volume and data frequency is critical to designing infrastructure for processing data and model stability [55]. To test the ability of stock selection of the model, three other factors ”QF_10%”, ”QF_5%”, ”QF_1%” are trained and tested (with ”QF_Month_1%, named as ”QF_20%”). The four groups are all trained under $d_{f}$ = 3, which include the first, middle, and tail 20%, 10%, 5%, and 1% of stocks in ranking, respectively. The backtest period and other settings are the same with 4.
Table 4: Result of factors under different scales
Strategy AR AER $\sigma_{p}$ WR SR Alpha VaR
QF_2017.35% 19.43% 0.162 57.8% 0.915 0.162 2.81%
QF_1013.12% 17.73 0.159 61.8% 0.574 0.128 2.023%
QF_512.59% 16.02% 0.136 61% 0.63 0.117 2.015%
QF_124.71% 35.74% 0.214 53.3% 0.967 0.249 3.048%
CSI300 1.77% \ 0.197 \ -0.015 \ 3.19%
Table 4 shows the result of the backtest. QF_10strategy recorded an AR of 13.12% and an AER of 17.73%, with a slightly lower portfolio volatility of 0.159. This strategy had the highest WR at 61.8%. QF_5strategy delivered 12.59% AR and 16.02% AW, with the lowest portfolio volatility at 0.136. It had the lowest VaR at 2.015%. QF_1strategy outperformed the others, achieving the highest AR of 24.71% and the highest AER of 35.74%, though it had the highest volatility at 0.214. These four factors all perform better than CSI300 based on their return and risk. The result shows that the model takes the ability to select stocks under different training set and requirements.
@@ 6 Conclusion
This study proposed a new model, quantformer for quantitative stock prediction and trading. We addressed the need for handling numerical input data rather than text and adapting the model for forecasting tasks rather than sequence-to-sequence problems common in NLP. To enable direct processing of numerical time series data, we replaced the word embedding layer with a standard linear layer and removed the output masking operations. We also simplified the decoder to produce a probability distribution over future price movements rather than autoregressively generating token sequences.
@@
There is further scope to enhance the framework, for instance by incorporating additional signals like news and fundamentals as auxiliary inputs. The self-attention mechanism could likely be improved to encode greater temporal context, or use transformer-based models such as GPT-4 [56] and Claude3 [57] to establish the strategy as these models are fine-tuned and may perform better for quantitative finance tasks.
Overall, our work illustrates the viability of quantformer for financial data modeling, provides a flexible framework to model the market from multiple perspectives of financial time series and develops profitable trading strategies that consider different variables. With further research to solve the above limitations, we believe such ML-based quantitative methods hold rich potential. The framework provides a flexible foundation for better understanding markets and developing profitable trading strategies.
The implementation code of quantformer is available at https://github.com/zhangmordred/QuantFormer.
@@ 7 Acknowledgements and declaration
This work was funded by the Natural Science Foundation of China (12271047); Guangdong Provincial Key Laboratory of Interdisciplinary Research and Application for Data Science, BNU-HKBU United International College (2022B1212010006); UIC research grant (R0400001-22; UICR0400008-21; R72021114); Guangdong College Enhancement and Innovation Program (2021ZDZX1046).
@@
Appendix A Detailed backtest results between 100 factors and the quantformer factor
All the data of factors comes from JoinQuant [^4] https://www.joinquant.com/, which is one of the most common quantitative finance platforms. Detailed descriptions of the following factors are available at JoinQuant Factors.
Ten metrics are used to evaluate the performance of each factor:
@@ Appendix A Detailed backtest results between 100 factors and the quantformer factor
  • Sharpe Ratio (SR) is a measure of risk-adjusted return that describes the additional earnings an investor receives for each standard deviation unit increase, which is shown in equation (9).
@@
  • The Alpha measures the ability of a portfolio to generate returns above the market benchmark. A positive alpha indicates that the investment portfolio has achieved a higher return than its benchmark after risk adjustment, which is shown in equation (10).
@@
  • Beta is the correlation between the portfolio and systemic risk, reflecting the sensitivity of the strategy to the change in the market. For the daily return of the strategy, $D_{p}$, and the daily return of the benchmark $D_{m}$, the equation of beta is:
    $Beta=\beta_{p}=\frac{Cov(D_{p},D_{m})}{Var(D_{m})}$
    where $Cov(D_{p},D_{m})$ is the correlation of between $D_{p}$ and $D_{m}$, $Var(D_{m})$ is the variance of daily return of benchmark.